### v1

# Covariance-Based Device Activity Detection with Massive MIMO for Near-Field Correlated Channels

Abstract
This paper studies the device activity detection problem in a massive multiple-input multiple-output (MIMO) system for near-field communications (NFC). In this system, active devices transmit their signature sequences to the base station (BS), which detects the active devices based on the received signal. In this paper, we model the near-field channels as correlated Rician fading channels and formulate the device activity detection problem as a maximum likelihood estimation (MLE) problem. Compared to the traditional uncorrelated channel model, the correlation of channels complicates both algorithm design and theoretical analysis of the MLE problem. On the algorithmic side, we propose two computationally efficient algorithms for solving the MLE problem: an exact coordinate descent (CD) algorithm and an inexact CD algorithm. The exact CD algorithm solves the one-dimensional optimization subproblem exactly using matrix eigenvalue decomposition and polynomial root-finding. By approximating the objective function appropriately, the inexact CD algorithm solves the one-dimensional optimization subproblem inexactly with lower complexity and more robust numerical performance. Additionally, we analyze the detection performance of the MLE problem under correlated channels by comparing it with the case of uncorrelated channels. The analysis shows that when the overall number of devices is large or the signature sequence length is small, the detection performance of MLE under correlated channels tends to be better than that under uncorrelated channels. Conversely, when is small or is large, MLE performs better under uncorrelated channels than under correlated ones. Simulation results demonstrate the computational efficiency of the proposed algorithms and verify the correctness of the analysis.
Index Terms:
Coordinate descent (CD) algorithm, device activity detection, massive random access, massive multiple-input multiple-output (MIMO), near-field communications (NFC).I Introduction
Massive machine-type communication (mMTC) plays a crucial role in fifth-generation (5G) and beyond cellular networks [1]. In mMTC, a key challenge is to handle massive random access, in which a large number of sporadically active devices are connected to the network [2]. This challenge can be addressed during the activity detection phase, in which each device is preassigned a unique nonorthogonal signature sequence, and the active devices transmit their signature sequences to the base station (BS). The BS then detects the active devices by identifying which sequences are transmitted based on the received signal [3].
There are generally two mathematical approaches for device activity detection. The first one is the compressed sensing (CS)-based approach, which jointly estimates the device activities and instantaneous channel state information (CSI) by exploiting the sporadic nature of device activities [4]. The second one is the covariance-based approach, which only estimates the device activities by taking advantage of the statistical CSI [5]. It formulates the activity detection problem as a maximum likelihood estimation (MLE) problem, which is referred to as the covariance-based approach because the MLE formulation only depends on the sample covariance matrix of the received signal. Studies have shown that the covariance-based approach not only outperforms the CS-based approach but also correctly detects a much larger number of active devices in massive multiple-input multiple-output (MIMO) systems [6, 7, 8].
The device activity detection problem has been extensively studied with respect to theoretical analysis, algorithm design, and practical issues, with most existing works focusing on classic far-field uncorrelated Rayleigh fading channels [6, 7, 8, 9, 10, 11, 12, 13]. However, with emerging trends such as extremely large-scale MIMO (XL-MIMO) and tremendously high frequencies in sixth-generation (6G) networks, the traditional far-field assumption is no longer accurate [14, 15]. This is because the Rayleigh distance, which separates the near-field and far-field regions, is proportional to the square of the antenna aperture and signal frequency. Consequently, many devices might be within the near-field region. Therefore, it is necessary to adopt the more practical near-field channel model for device activity detection.
In this paper, we study the covariance-based activity detection problem under the spatial correlated Rician fading channel, where the deterministic line-of-sight (LoS) components and correlation matrices are derived from near-field spherical wave propagation [14, 15, 16, 17]. This setup makes both algorithm design and detection performance analysis more challenging than those under traditional uncorrelated Rayleigh fading channels, primarily due to the channel correlation. For spatially uncorrelated channels, the MLE formulation relies on the marginal distributions of each column of the received signal, as these columns are independently distributed [6]. However, in the case of spatially correlated channels, different columns of the received signal are dependent with each other and it is necessary to consider their joint distribution. The joint distribution complicates the MLE problem because its dimension is much higher than that of the marginal distributions and increases with the number of antennas. Moreover, since the existing detection performance analyses [6, 7, 8] assume that different columns of the received signal are independent and identically distributed (i.i.d.), they cannot be directly applied to the case of spatially correlated channels.
I-A Related Works
The CS-based approach is an important line of research in device activity detection under uncorrelated Rayleigh fading channels, which takes advantage of the sporadic nature of the device activity pattern. Existing sparse recovery methods for solving the CS-based activity detection problem include approximate message passing (AMP) [18, 19, 20, 21, 22] and group least absolute shrinkage and selection operator (LASSO) [23, 24, 25].
Another important line of research in device activity detection is the covariance-based approach, which exploits the statistical information of the channel and formulates the activity detection problem as an MLE problem. From an algorithmic perspective, the coordinate descent (CD) algorithm [5] and its accelerated variants [26, 8, 27] are the most popular and efficient algorithms for solving the MLE problem. Other algorithms include the expectation maximization/minimization (EM), also known as sparse Bayesian learning [28], the sparse iterative covariance-based estimation (SPICE) [29], and gradient descent [30]. From a theoretical perspective, the covariance-based approach has been demonstrated to have a quadratic scaling law, which enables it to detect many more active devices than the CS-based approach [6, 7, 8]. However, the above algorithms and theories all rely on the assumption of uncorrelated Rayleigh fading channels and are not directly applicable to correlated Rayleigh/Rician fading channels.
The CS-based approach under correlated Rayleigh fading channels has been studied in [31, 32, 33, 34, 35, 36, 37, 38]. Specifically, by utilizing the channel distribution information, the AMP algorithm [31, 32, 33] and orthogonal AMP (OAMP) algorithm [39, 34] are proposed for solving the device activity detection problem. The works [35, 36] study the problem formulation of the CS-based activity detection and utilize the alternating direction method of multipliers (ADMM) algorithm for solving it. Additionally, [37] proposes a Bayesian framework for CS-based activity detection with unknown correlation matrices, and [38] proposes a turbo-type compressive sampling matching pursuit (Turbo-CoSaMP) algorithm for solving the device activity problem by exploiting the structured sparsity of the near-field channel.
However, there are few studies on the covariance-based approach under correlated Rayleigh or uncorrelated Rician fading channels. Three works [40, 41, 42] investigate the covariance-based approach under correlated Rayleigh fading channels. In [40], the MLE formulation and algorithm are studied under the assumption that the eigenvectors for the correlation matrices of all devices are identical. In [41], the covariance-based approach is applied to intelligent reflecting surface (IRS)-aided activity detection, and a deep unfolding approach is proposed for solving the MLE problem. The work [42] studies the covariance-based approach when the channels are correlated within time-frequency blocks. Moreover, under uncorrelated Rician fading channels, the problem formulation and algorithms of the covariance-based approach are studied in [43] and [44]. In this paper, we study the algorithm design and theoretical analysis of the covariance-based approach considering correlated Rician fading channels.
The covariance-based approach for device activity detection has been applied in various practical scenarios, including the joint activity and data detection problem [9, 45, 46], the asynchronous system with joint activity and delay detection [10, 11], the multi-cell system [8, 47], the cell-free system [48, 49], the unsourced random access scenario [6], the system with frequency offsets [50, 43], the system with low-resolution analog-to-digital converters [51], and the orthogonal frequency division multiplexing (OFDM)-based massive grant-free access scheme [52]. Moreover, the deep learning approach is applied to device activity detection [53]. Finally, the covariance-based approach can be used in a three-phase protocol [54, 55] to first identify active devices and then schedule them in orthogonal transmission slots for channel estimation.
I-B Main Contributions
This paper investigates the covariance-based activity detection problem in massive MIMO systems under spatially correlated Rician fading channels. The device activity detection problem is formulated as an MLE problem. The correlation of channels complicates both the algorithm design and theoretical analysis of the MLE problem. First, due to the channel correlation, when applying the CD algorithm, there is no closed-form solution to the one-dimensional optimization subproblem, and the dimension of the covariance matrix in the MLE formulation is proportional to the number of antennas, which is significantly different from the case where the channels are uncorrelated. Second, compared with the uncorrelated channel case, the channel correlation makes the columns of the received signal no longer i.i.d., which challenges the theoretical analysis of MLE’s detection performance. The main contributions of the paper are twofold.
On the algorithmic side, we propose two computationally efficient algorithms for solving the MLE problem: an exact CD algorithm and an inexact CD algorithm. The exact CD algorithm iteratively solves the one-dimensional optimization subproblem exactly using matrix eigenvalue decomposition and polynomial root-finding. However, due to its high complexity and potential numerical errors when the rank of the channel correlation matrix is large, the exact CD algorithm is not efficient and sometimes is even divergent. To address this issue, we propose the inexact CD algorithm, which primarily relies on Taylor expansion to approximate some terms in the objective function. The resulting inexact CD algorithm has a lower complexity and is guaranteed to converge. Additionally, we study a special case where only a few devices’ channels are correlated with a low-rank correlation matrix, while the channels of many other devices are uncorrelated. We demonstrate that exploiting the low-rank properties of the correlation matrices significantly reduces the complexity of both the exact CD and inexact CD algorithms.
On the analytical side, we analyze the detection performance of the MLE problem under correlated channels by comparing it with the case of uncorrelated channels. This comparison is examined from two perspectives: signal statistical dimension and identifiability of the true activity. From the signal statistical dimension perspective, due to the low-rank property of the near-field spatial correlation matrix, the detection performance of MLE under correlated channels tends to be worse than that under uncorrelated channels when the overall number of devices is small or the signature sequence length is large. Conversely, from the perspective of identifiability, correlated channels make the identifiability of the true activity more likely to hold than uncorrelated channels, suggesting that MLE performs better in correlated channels than in uncorrelated ones when is large or is small.
I-C Paper Organization and Notations
The rest of the paper is organized as follows. Section II introduces the system model and the MLE formulation. Section III presents the efficient CD algorithms for solving the MLE problem. In Section IV, we analyze the difference in detection performance between correlated and uncorrelated channels. Simulation results are provided in Section V. Finally, we conclude the paper in Section VI.
In this paper, lower-case, boldface lower-case, and boldface upper-case letters denote scalars, vectors, and matrices, respectively. Calligraphic letters denote sets. Superscripts and denote conjugate transpose, transpose, conjugate, and inverse, respectively. Additionally, denotes the imaginary unit, denotes the identity matrix of dimension denotes expectation, denotes the trace of (or ) denotes a (block) diagonal matrix formed by (or ), denotes definition, denotes the determinant of a matrix, or the (coordinate-wise) absolute operator, denotes the -norm of and denotes the Kronecker product. Finally, denotes a complex Gaussian distribution with mean and covariance
II System Model and Problem Formulation
II-A System Model
We consider an uplink single-cell massive MIMO system consisting of a BS equipped with an -antenna uniform linear array with half-wavelength antenna spacing. In the cell, there are single-antenna devices, and only devices are active during any coherence interval due to the sporadic traffic. We use to denote the binary variable with for active and for inactive devices. We consider the case where the devices are in the near-field region, i.e., the distance between the devices and the BS is less than the Rayleigh distance, which is given by where and denote the antenna aperture and the wavelength, respectively [14]. We assume that both the deterministic LoS channel and statistical multipath channels exist, and the channel between device and the BS can be modeled as the following Rician fading model [14, Eq. (133)]:
where is the LoS channel between device and the BS, and
where denotes the channel gain between device and the BS, is the distance between the device and the -th antenna, and is the distance between the device and the midpoint of the antenna array. The correlation matrix is given by [14, Eq. (134)]:
where is the number of scatterers between device and the BS, is the intensity attenuation caused by the -th scatterer, denotes the channel gain between the -th scatterer and the device and is the channel vector between the -th scatterer and the BS with a similar structure to (2), i.e., where and denote the corresponding channel gain and distances. Notice that and thus is a low-rank matrix when the number of scatterers
Each device is preassigned a unique signature sequence with length for activity detection. During the uplink pilot stage, all active devices synchronously transmit their signature sequences to the BS as random access requests. Following this, the received signal at the BS can be expressed as
where is the additive i.i.d. white Gaussian noise with variance Denote as the activity indicator vector.
II-B Problem Formulation
The task of device activity detection is to estimate from the received signal Notice that the Rician fading components and noises are both Gaussian. Therefore, the received signal in (4) follows a Gaussian distribution. Since the columns of are generally not i.i.d., we need to consider the joint distribution of all its columns, i.e., the distribution of its vectorization:
where For a given is a Gaussian vector, i.e., where the mean is
and the covariance matrix is given by
Therefore, the likelihood function of can be expressed as
The MLE problem can be formulated as the minimization of the negative log-likelihood function which can be expressed as
Problem (9) is computationally intractable due to its binary constraints (9b). To make the problem more computationally tractable, we relax the constraint to as in [6, 7, 8], and obtain the following relaxed problem:
In this paper, we focus on the algorithm design and performance analysis for the MLE problem (10). Notice that the dimension of the covariance matrix in problem (10) is which is significantly different from the case of uncorrelated channels. In the latter case, the covariance matrix involved in the MLE formulation is [5, 6, 7]. This difference makes problem (10) much more challenging to solve.
III Proposed CD Algorithms for Device Activity Detection
In this section, we focus on designing efficient algorithms for solving the MLE problem (10). Notice that problem (10) is nonconvex due to the concavity of We propose two algorithms for solving the MLE problem: an exact CD and an inexact CD. In the exact CD algorithm, we solve the one-dimensional optimization subproblem exactly, while in the inexact CD algorithm, we solve it inexactly in order to reduce the computational cost. Furthermore, we study a special case that allows us to further reduce the complexity of both the exact CD and inexact CD algorithms by utilizing the low-rank property of the near-field channel correlation matrix.
We first introduce several key symbols and present the mathematical formulations used to solve the MLE problem. Let Then there exists a matrix such that
Let Then we have
Let denote the objective function of problem (10). The -th component of the gradient of is given by
Please refer to Appendix A for the derivation of (13). Next, we present the first-order optimality condition of problem (10):
which is a necessary condition for the optimality of To quantify the degree to which each coordinate of violates condition (14), we define a nonnegative vector as follows:
where is the (coordinate-wise) projection operator onto the interval and is the (coordinate-wise) absolute value. Notice that the first-order optimality condition (14) is equivalent to Our goal is to find a feasible point such that where is a given error tolerance.
III-A Exact and Inexact CD Algorithms
We use the framework of randomly permuted CD [9], which is one of the most efficient variants of the CD algorithm. At each iteration, the algorithm randomly permutes the indices of all coordinates and updates them sequentially according to this permutation. Let denote the vector whose -th component is and the other components are For any given coordinate the algorithm updates by solving the following one-dimensional optimization subproblem:
Notice that problem (16) generally does not have a closed-form solution, which sharply contrasts with the CD algorithm in the case where the channels are uncorrelated [5, 6, 7].
The key to the CD algorithm is the efficient solution of problem (16). Next we analyze the property of its objective function. According to the properties of the determinant, we have
Furthermore, applying the Sherman-Morrison-Woodbury formula, we obtain
Substituting (III-A) and (18) into (16), we can rewrite problem (16) as (19), shown at the top of the next page.
Problem (19) is computationally more tractable than problem (16) for the following reason. The key to solving problem (19) is analyzing the determinant and inverse of the low-dimensional matrix In contrast, directly solving problem (16) involves handling the high-dimensional matrix where
We present two methods for solving problem (19) below. Once it is solved, we update variables and with being updated according to (18). The exact CD and inexact CD algorithms are summarized in Algorithm 1.
III-A1 Solving Problem (19) Exactly
In this part, we present a method for solving problem (19) exactly, which performs well when is small. The main computational complexities are matrix eigenvalue decomposition and polynomial root-finding.
We now reformulate problem (19) by addressing each term separately. First, we consider the term in (19a). According to the eigenvalue decomposition of the matrix there exists a unitary matrix and a diagonal matrix that satisfy
Then, we have
which is a polynomial of degree in Second, the term in (19b) is a tractable quadratic polynomial. Next, we consider the terms in (19c)–(19e). These terms have the same structure: with For example, in (19e), and We next analyze Applying (20) and setting and we get
which is a fraction where the denominator is a polynomial of degree and the numerator is a polynomial of degree Notice that is also the polynomial in (21). By substituting different and into (III-A1), we can compute the coefficients of the numerator of each term in (19c)–(19e). We then combine all terms in (19b)–(19e) and reduce them to common denominator Finally, we rewrite problem (19) as
where is a polynomial of degree and its coefficients are derived by combining all the terms in (19b)–(19e).
The process for solving problem (23) is as follows. First, we find such that the derivative of the objective function in (23) is zero. This is equivalent to setting the numerator of the derivative to zero, which gives:
where and are the derivatives of and respectively. Notice that the left-hand side in (24) is a polynomial of degree so problem (24) can be solved using a polynomial root-finding algorithm [56, Chap. 6]. Next, we choose the solution with the smallest objective function value in (23) among the roots of (24) and the two boundary points and This yields the exact solution to problem (23) (i.e., problem (19)). The procedure of solving problem (19) exactly is outlined in lines 6 and 7 of Algorithm 1.
The above proposed method is suitable to solve problem (19) for small due to the following reasons. When is large, significant numerical errors arise in two steps: (i) the eigenvalue decomposition of i.e., the computation of and satisfying (20), and (ii) the root-finding of the polynomial of degree in (24). The error in step (i) makes the coefficients of and inaccurate, while the error in step (ii) makes it impossible to obtain a solution to problem (23). To address these issues, we propose a method for solving problem (19) inexactly, as detailed below.
III-A2 Solving Problem (19) Inexactly
The key idea for solving problem (19) inexactly is to use the Taylor expansion to approximate the objective function in (19), thereby avoiding the eigenvalue decomposition and polynomial root-finding. Consequently, this method performs well when is large.
As follows, we approximate each term in (19). First, using (21) and the Taylor expansion, we have
where the little- notation satisfies the condition that as Further, we apply the Taylor expansion to obtain
Substituting (III-A2) and (III-A2) into (19) and ignoring we obtain an approximation of the objective function in (19), shown in (III-A2) at the top of the next page,
which is a quartic polynomial. Notice that calculating the coefficients in (III-A2) does not involve the eigenvalue decomposition. Next, we solve the following problem:
where is a properly selected parameter. The complexity of solving problem (28) is because all roots of the derivative of (28) can be obtained using the cubic formula. The optimal solution to (28) can be obtained by selecting the one with the smallest objective value among the roots and the two boundary points and The process of solving problem (19) inexactly is summarized in lines 9 and 10 of Algorithm 1.
The following theorem guarantees that, for an appropriate the solution to problem (28) sufficiently decreases the objective function of problem (19).
Theorem 1
Proof:
Please see Appendix A. ∎
Theorem 1 demonstrates that, at each coordinate update, although the inexact CD algorithm does not solve problem (19) exactly, it finds a good approximate solution that sufficiently decreases the objective function, especially when is large. This guarantees the convergence of the inexact CD algorithm (to a stationary point).
III-A3 Complexity
In Algorithm 1, the dominant computational complexity of updating a coordinate includes: (i) solving the one-dimensional problem (19), and (ii) updating matrix with complexity The complexity of solving problem (19) exactly is which includes the eigenvalue decomposition of matrix with complexity and the root-finding of the polynomial of degree with complexity The complexity of solving problem (19) inexactly is Therefore, the overall complexity of updating is
In massive MIMO systems, updating and storing dominate the complexity of both the exact and inexact CD algorithms. As follows, we consider a special case where the complexity of updating and storing can be significantly reduced.
III-B A Special Case Study
In this subsection, we consider a special case in which the proposed exact CD and inexact CD can be accelerated. The motivation arises from the fact that within a cell, there may be only a few devices in the near-field region of the BS, whereas the majority of devices are located in the far-field region. Furthermore, we assume that for each device in the near-field region, the number of scatterers is small. In this scenario, we can model the near-field channel as correlated channel with low-rank correlation matrix (since ), and assume that the far-field channel is uncorrelated. In this scenario, we only focus on reducing the complexity of updating and storing by taking advantage of the low-rank properties of the near-field correlation matrices. Below, we formally present the assumption for this specific case.
Assumption 1
Assume that there exists such that for the correlation matrix of device is a low-rank positive semidefinite matrix, and for the correlation matrix is where is the large-scale fading coefficient. We further assume that the summation of the correlation matrices of the first devices is rank-deficient, i.e.,
Notice that Then, a sufficient condition for Assumption 1 to hold is that there are few near-field devices (i.e., is small) and few scatterers in the near-field channels (i.e., is small). The following proposition demonstrates that under Assumption 1, () can be transformed into a matrix where only the upper-left submatrix is nonzero.
Proposition 1
Under Assumption 1, there exists a unitary matrix such that
where The matrix can be computed by performing the eigenvalue decomposition on the matrix i.e., finding a unitary matrix such that
where
Proof:
Please see Appendix B. ∎
Furthermore, once the matrix is obtained, we can transform the MLE problem (10) to obtain a more tractable problem as presented in the following theorem.
Theorem 2
Proof:
Please see Appendix C. ∎
An intuitive explanation of Theorem 2 is as follows. It is easy to verify that problem (33) corresponds to minimizing the negative log-likelihood function of in (35), i.e., According to the property of the transformation of the distribution function, we have
Therefore, problem (33) (i.e., minimizing ) is equivalent to problem (10) (i.e., minimizing ).
Notice that problem (33) has a similar structure to problem (10). Therefore, we can apply both exact and inexact CD algorithms in Algorithm 1 to solve it. In line 14 of Algorithm 1, we need to update the matrix with complexity However, can be updated and stored with significantly lower complexity. Specifically, in (2) is a block diagonal matrix, i.e.,
where
and is because we set for where satisfies and for Therefore, in a coordinate update, we can obtain the updated by updating two low-dimensional matrices and The details are summarized in Algorithm 2. Notice that the complexity of updating in Algorithm 2 is for and for which is significantly lower than because Furthermore, we only need to store and instead of because all matrix-vector multiplications involving can be expressed using and Specifically, for any vectors we have
Under Assumption 1, Algorithm 2 reduces the complexity of updating from to at most which is independent of Therefore, it will significantly reduce the complexity of exact and inexact CD algorithms in massive MIMO systems. The one-dimensional optimization subproblem can be solved either exactly or inexactly using techniques similar to those in Algorithm 1, with the same complexity. We omit these details here for simplicity.
If the condition (30) in Assumption 1 does not hold, i.e., we can design a low-complexity algorithm inspired by Algorithm 2 to obtain an approximate solution to the MLE problem (10). We describe it briefly as follows: First, we perform the eigenvalue decomposition on the matrix to obtain the unitary matrix Although the matrix does not retain the property in (31), we can select a parameter such that and set to the upper-left submatrix of i.e., for Next, we substitute into (2) and apply the exact or inexact CD algorithm to solve problem (33), with updated by Algorithm 2. The complexity of updating is then at most For a properly selected the output of the algorithm is a high-quality approximate solution to problem (10). This approximate solution can be used as an initial point for Algorithm 1 to accelerate the convergence of the algorithm.
IV Detection Performance Analysis
In this section, we analyze how the channel correlation matrices affect the detection performance of the MLE model (10). Our analysis is based on the low-rank properties of near-field spatial correlation matrices [14, 15, 16, 17], focusing on two perspectives: signal statistical dimension and identifiability of the true activity vector. On the one hand, low-rank correlation matrices may reduce the signal statistical dimension compared to uncorrelated channels, especially when the ratio is small. This reduction of signal statistical dimension will lead to a loss in detection performance. On the other hand, covariance identifiability under correlated channels is more likely to hold, implying that detection performance under correlated channels is better than that under uncorrelated channels when the ratio is large.
In order to reveal how the channel correlation affects the detection performance of the MLE model, the analysis in this section focuses on the following two cases, which differ only in the correlation of the channels:
- Case I:
Case II:
For each device , the channel follows where is the large-scale fading coefficient.
We assume that the set of signature sequences is the same in both Case I and Case II, with each sequence being randomly generated from a specific distribution, such as the uniform distribution over the discrete set as in [8].
IV-A Signal Statistical Dimension
First we define the signal statistical dimensions and in Case I and Case II respectively:
The statistical dimensions and are closely related to the detection performance in Case I and Case II due to the following reason. Without loss of generality, we consider Case I. Notice that the received signal follows where and are given in (6) and (II-B). Using techniques similar to Proposition 1, we have that there exists a unitary matrix such that
where Defining the transformed received signal we have where and
We observe from (45) that is a block diagonal matrix and its lower-right submatrix is independent of Therefore, in the transformed received signal only the covariance of subvector depends on while the covariance of subvector is just Applying the property of the transformation of the distribution function, we can rewrite the log-likelihood function as
where is due to the fact that and are independent of each other for a given The log-likelihood function in (IV-A) implies that when solving the MLE problem (10), only the mean and covariance of subvector and the mean of subvector are used for activity detection. This is why we refer to as the statistical dimension. Furthermore, if for all then subvector is simply the noise that follows In this case, only the -dimensional vector is useful for activity detection. Therefore, the larger the value of the better the detection performance in Case I. The same relationship holds for Case II with respect to
Since and in (42) and (43) generally do not have a closed-form expression, we estimate their values by calculating their upper bounds. In Case I, denote Then we obtain
Since also satisfies we have as an upper bound on In Case II, due to and using similar techniques as in (IV-A), we have the upper bound where is due to the fact that in the random access setting Therefore, we have
The key inference from (48) is that when the ratio is small, the detection performance in Case I would be worse than that in Case II. Whereas when is large, and are close to the same. Next, we will illustrate from another perspective that the detection performance in Case I is better when is large.
IV-B Identifiability of True Activity Vector
The phase transition analysis in [7] demonstrates that, a crucial condition for the MLE model (10) to perform effectively is that the true activity indicator vector needs to be uniquely identifiable. This implies that there should be no other vector such that We now investigate the identifiability of in both Case I and Case II. Since both and are multivariate Gaussian distributions, their equivalence implies that their corresponding means and covariances must be identical. Therefore, the identifiability of is equivalent to the condition that there is no other such that and Furthermore, we define
The identifiability of in Case I and Case II is equivalent to the conditions and respectively. The following theorem illustrates that is a sufficient condition for
Theorem 3
If then also holds true.
Proof:
Please see Appendix D. ∎
Theorem 3 highlights a key observation that the identifiability of is more likely to hold in Case I than in Case II. Furthermore, the scaling law results in [6, 7, 8] show that when the system parameters satisfy and under certain methods of generating signature sequences, holds. Therefore, also holds under these conditions, implying that the number of active devices that can be correctly detected by the MLE problem (10) in Case I is at least
Notice that if for all then sets and can be represented as
where and are matrices whose -th columns are and respectively. Taking Case I as an example, an heuristic view is that if the columns of are closer to orthogonality, then the condition is more likely to hold, which also implies that the mutual interference between devices is smaller. To this end, for different devices and we study the following cosine similarity:
where follows from the Cauchy-Schwartz inequality and equality holds if and only if there exists a constant such that We observe from (54) that and are closer to orthogonal than and implying that the mutual interference between devices and in Case I is lower than that in Case II. The above analysis demonstrates that the channel correlation among different antennas can reduce the correlation between devices and thereby improve the detection performance. Therefore, when the signal statistical dimensions in Case I and Case II are equal (i.e., is large), the MLE model (10) exhibits superior detection performance in Case I (than in Case II).
V Simulation Results
We consider an mMTC system with a single cell of radius m, where all potential devices are uniformly distributed. In the simulation, the carrier frequency is set as GHz. The scatterers are uniformly distributed within m from the BS, and the intensity attenuation is for all The channel gain and follow the path-loss model, which is given by where is the corresponding distance in km. We adopt a power control scheme such that where the power is set as dBm. The background noise power is set as dBm/Hz over MHz. The signature sequences are generated randomly and uniformly from the discrete set as in [8].
V-A Computational Efficiency of Proposed Algorithms
In this subsection, we showcase the efficiency of the proposed exact and inexact CD algorithms for solving the device activity detection problem (10). In the inexact CD algorithm, we set the parameter to
In Fig. 1, we show the decrease of the objective function values of the proposed exact and inexact CD algorithms with respect to the running time with and The outputs of the exact and inexact CD algorithms are denoted as and respectively. We then define as the minimum of and We observe from Fig. 1(a) that when the objective function values of exact and inexact CD can gradually decrease and converge to the same value. Moreover, exact CD is significantly more efficient than inexact CD. This is because, when is small (implying is also small), the exact CD algorithm can find the exact solution to the one-dimensional optimization subproblem (19) in each coordinate update, whereas the inexact CD algorithm can only find an approximate solution. In Fig. 1(b), we observe that when the objective function value of the inexact CD algorithm still converges, but that of the exact CD algorithm cannot decrease and diverges. This occurs because when is large (corresponding to a large ), the exact CD algorithm fails to solve the one-dimensional subproblem (19) due to the numerical errors associated with the high-dimensional eigenvalue decomposition and polynomial root-finding. This observation shows that exact CD is more efficient when is small, while inexact CD is more robust when is large.
In Fig. 2, we assume that Assumption 1 holds true and plot the decrease of the objective function values of the proposed exact and inexact CD algorithms as well as their accelerated versions by Algorithm 2 with respect to the running time with and In Fig. 2(a), we set in Assumption 1, and for (implying that ). Consequently, we have in (30). Notice that in this case, there are devices with channel correlation matrix Fig. 2(a) shows that, in this scenario, the objective function value of the accelerated inexact CD converges to and it is more computationally efficient than the inexact CD algorithm. However, neither exact CD nor its accelerated version converges. This is because neither exact CD nor its accelerated version can solve the one-dimensional optimization subproblem (19) when the channel correlation matrix of the device To demonstrate that the exact CD algorithm can also be accelerated when it converges, we remove devices with channel correlation matrix Therefore, in Fig. 2(b), we set in Assumption 1, and for all We observe from Fig. 2(b) that the accelerated version of exact CD converges, and it is more efficient than exact CD. The same observation applies to the inexact CD algorithm as well.
A summary of the proposed CD algorithms is as follows. First, the exact CD algorithm is more computationally efficient than the inexact CD algorithm when the rank of the channel correlation matrix is small. Second, the inexact CD algorithm is more favorable in terms of robustness in the case where the rank of the channel correlation matrix is large. Finally, both of the exact and inexact CD algorithms can be accelerated when Assumption 1 holds. Due to its robustness in solving the subproblem, we adopt the inexact CD algorithm to solve problem (10) in different scenarios in the following two subsections.
V-B Numerical Validation of Detection Performance Analysis
In this subsection, we verify the detection performance analysis in Section IV. We consider the device activity detection problem under Case I and Case II and perform MLE model (10) in both cases. In order to demonstrate the impact of channel correlation matrices on the detection performance as compared to the traditional uncorrelated Rayleigh fading channels, we set for all for a fair comparison.
In Fig. 3, we show the detection performance for different values of with and The “probability of error” in the figure is defined as the point where the probability of missed detection (PM) and probability of false alarm (PF) are equal. We observe that increasing can improve the detection performance in both cases. When is small, the detection performance in Case I is better, and when is large, the detection performance in Case II is better. This is because is rank-deficient in Case I and full-rank in Case II, which makes the signal statistical dimension (defined in (42) and (43)) in Case II larger than that in Case I when is large. In other words, in Case I, some dimensions of the received signal are actually useless. We can also observe from Fig. 3 that in Case I, when the larger is, the better the detection performance is. This also verifies when is larger, the rank of is larger, and thus the signal statistical dimension is larger, and the detection performance is better.
Fig. 4 plots the detection performance for various values of with parameters and It is observed that the detection performance of both cases deteriorates as increases. Furthermore, when is large, for instance, the detection performance in Case I is superior to that of Case II. This is due to when is large, the signal statistical dimensions are equal in both cases; however, in Case I, the true activity vector is more easily identifiable compared to Case II, as discussed in Section IV-B. Additionally, we also observe that in Case I, a smaller tends to result in better detection performance for larger This can be explained by the fact that when is smaller, the cosine similarity in (54) is likely to be lower, indicating a weaker correlation between different devices, which in turn leads to improved detection performance.
Fig. 5 shows the detection performance for various values of with and We observe that as increases, the detection performance in Case I with values of and significantly outperforms that of Case I with and Case II. This suggests that when is large and the rank of is large in Case I, the detection performance is superior compared to the case when the rank of is low in Case I, and it is also superior to Case II.
V-C Detection Performance Comparison with Benchmarks
In this subsection, we compare the detection performance of the proposed MLE-based approach with the following benchmarks:
AMP [31] and OAMP [34]: Both algorithms are designed for joint activity detection and channel estimation under correlated Rayleigh fading channels. OAMP is an improved version of AMP, which is applicable to more types of signature sequences. In the scenario of Rician fading channels, i.e., we ignore the LoS channel and assume the prior distribution to be in both methods. Notice that both methods require the knowledge of the probability that a device is active, i.e., but the MLE does not require this prior information.
Mismatched MLE: We apply the vanilla MLE-based approach in [43]. Specifically, we use to replace the channel correlation matrix in the MLE formulation.
In Fig. 6, we show the detection performance under correlated Rician/Rayleigh fading channels for various values of with and It can be observed from Fig. 6 that increasing can improve the detection performance of all compared methods. In Fig. 6(a), we observe that the detection performance of MLE is significantly better than that of all the other benchmarks. The main reason for this is that these benchmarks use inaccurate channel distributions: both AMP and OAMP do not utilize the LoS channel, and the mismatched MLE roughly approximates In Fig. 6(b), we observe that the detection performance of MLE is slightly better than that of AMP and OAMP when is large and significantly better than that of mismatched MLE. In this case, all methods except mismatched MLE use accurate channel distributions. Although the detection performance of MLE, AMP, and OAMP is relatively close, a significant advantage of MLE is that it does not require the prior knowledge of
Fig. 7 plots the probability of error for different values of with and We observe that increasing can improve the detection performance of all compared methods. In Fig. 7(a), due to the use of inaccurate channel distributions, the detection performance of AMP, OAMP, and mismatched MLE is significantly worse than that of MLE. Moreover, when is large, the detection performance of mismatched MLE becomes better than that of AMP and OAMP. In Fig. 7(b), we observe that the detection performances of MLE, AMP, and OAMP are very close, and the detection performance of mismatched MLE is significantly worse than theirs.
Fig. 8 plots the probability of error for various values of with and We observe that as increases, the detection performance of the compared methods deteriorates. Fig. 8(a) illustrates that in the Rician fading scenario, the detection performance of MLE is significantly better than that of the other methods. Fig. 8(b) shows that in the Rayleigh fading scenario, the detection performance of MLE, AMP, and OAMP is very close, with the mismatched MLE performing worse than these methods. Moreover, MLE outperforms AMP and OAMP when is large.
VI Conclusion
This paper investigates the device activity detection problem for NFC, where the near-field channel can be modeled as correlated Rician fading. The corresponding device activity detection problem is formulated as an MLE problem. Two efficient algorithms for solving the MLE problem, i.e., exact and inexact CD, are proposed. In addition, an analysis of the detection performance of MLE is provided. Simulation results demonstrate the efficiency of the proposed algorithms and verify the correctness of the detection performance analysis.
Appendix A Proof of Theorem 1
The outline of the proof is as follows. First, we prove that for a sufficiently large is an upper bound of We then show that inequality (29) holds by studying the upper bound in more detail.
According to (III-A2) and (III-A2), we have
The gradient in (13) is derived from (III-A2) and
Since is continuous and is in the bounded region is Lipschitz continuous. That is, there exists a constant such that for any the following inequality holds:
Furthermore, we have a quadratic upper bound on [57, Proposition A.24]:
Setting and in (58), we obtain
Next, we combine all the homogeneous terms in (III-A2) and denote as the coefficient of Specifically,
Since and are continuous and is in the bounded region they are bounded. That is, there exist constants and such that
Therefore, setting for we have that for any the following inequality holds
Combining (59) and (62), we obtain
Now we study the optimal value of to obtain the desired inequality (29). Notice that for any
Therefore, the optimal value of satisfies
where is derived from
and the properties of the projection operator Similar techniques are used in the proof of [8, Proposition 4]. Substituting into (63) and combining it with (A), we obtain (29). This completes the proof of Theorem 1.
Appendix B Proof of Proposition 1
Appendix C Proof of Theorem 2
Appendix D Proof of Theorem 3

### v2

@@
 Compared to the traditional uncorrelated channel model, the correlation of channels complicates both algorithm design and theoretical analysis of the MLE problem.
-On the algorithmic side, we propose two computationally efficient algorithms for solving the MLE problem: an exact coordinate descent (CD) algorithm and an inexact CD algorithm.
-The exact CD algorithm solves the one-dimensional optimization subproblem exactly using matrix eigenvalue decomposition and polynomial root-finding.
-By approximating the objective function appropriately, the inexact CD algorithm solves the one-dimensional optimization subproblem inexactly with lower complexity and more robust numerical performance.
+On the algorithmic side, we present the classical exact coordinate descent (CD) algorithm for solving the MLE problem, which suffers from numerical instability when applied to correlated channels.
+We propose a computationally efficient inexact CD algorithm by approximating the objective function, which approximately solves the one-dimensional subproblem and improves both computational efficiency and numerical stability.
 Additionally, we analyze the detection performance of the MLE problem under correlated channels by comparing it with the case of uncorrelated channels.
@@
 Conversely, when is small or is large, MLE performs better under uncorrelated channels than under correlated ones.
-Simulation results demonstrate the computational efficiency of the proposed algorithms and verify the correctness of the analysis.
+Finally, we study the MLE model in the joint device activity and data detection context.
+Simulation results demonstrate the computational performance of the presented algorithms and verify the correctness of the analysis.
 Index Terms:
@@
 Therefore, it is necessary to adopt the more practical near-field channel model for device activity detection.
-In this paper, we study the covariance-based activity detection problem under the spatial correlated Rician fading channel, where the deterministic line-of-sight (LoS) components and correlation matrices are derived from near-field spherical wave propagation [14, 15, 16, 17].
+In this paper, we study the covariance-based activity detection problem under the spatially correlated Rician fading channel, where the deterministic line-of-sight (LoS) components and correlation matrices for the near-field devices are derived from near-field spherical wave propagation [15, 16, 17, 18].
 This setup makes both algorithm design and detection performance analysis more challenging than those under traditional uncorrelated Rayleigh fading channels, primarily due to the channel correlation.
@@
 However, in the case of spatially correlated channels, different columns of the received signal are dependent with each other and it is necessary to consider their joint distribution.
-The joint distribution complicates the MLE problem because its dimension is much higher than that of the marginal distributions and increases with the number of antennas.
+The joint distribution complicates the MLE problem because its dimension is which is much larger than that of the marginal distribution (i.e., ), where is the signature sequence length and is the number of antennas.
+This is particularly relevant in the context of massive MIMO.
 Moreover, since the existing detection performance analyses [7, 8, 9] assume that different columns of the received signal are independent and identically distributed (i.i.d.), they cannot be directly applied to the case of spatially correlated channels.
@@
 The main contributions of the paper are twofold.
-On the algorithmic side, we propose two computationally efficient algorithms for solving the MLE problem: an exact CD algorithm and an inexact CD algorithm.
-The exact CD algorithm iteratively solves the one-dimensional optimization subproblem exactly using matrix eigenvalue decomposition and polynomial root-finding.
-However, due to its high complexity and potential numerical errors when the rank of the channel correlation matrix is large, the exact CD algorithm is not efficient and sometimes is even divergent.
-To address this issue, we propose the inexact CD algorithm, which primarily relies on Taylor expansion to approximate some terms in the objective function.
-The resulting inexact CD algorithm has a lower complexity and is guaranteed to converge.
+On the algorithmic side, we introduce an exact CD algorithm for solving the MLE problem, which is a direct extension of the CD algorithm designed for the uncorrelated channel case [6].
+It exactly solves the one-dimensional optimization subproblem by using matrix eigenvalue decomposition and polynomial root-finding techniques.
+However, this algorithm suffers from numerical instability and can fail to converge when the channel correlation matrix has a high rank.
+To address this issue, we propose an inexact CD algorithm, which primarily relies on Taylor expansion to approximate selected terms in the objective function.
+This approach provides improved numerical stability and reduced per-coordinate update complexity.
 Additionally, we study a special case where only a few devices’ channels are correlated with a low-rank correlation matrix, while the channels of many other devices are uncorrelated.
@@
 Conversely, from the perspective of identifiability, correlated channels make the identifiability of the true activity more likely to hold than uncorrelated channels, suggesting that MLE performs better in correlated channels than in uncorrelated ones when is large or is small.
+In addition, we study the joint device activity and data detection problem, where each active device wishes to send a few bits of information to the BS.
+In this scenario, each device is associated with multiple signature sequences and can embed a small number of information bits.
+We propose the MLE model for this scenario.
+Simulation results show that transmitting several bits of information does not significantly deteriorate the detection performance of the MLE model, especially when the length of the signature sequence is large.
 I-C Paper Organization and Notations
@@
 In Section IV, we analyze the difference in detection performance between correlated and uncorrelated channels.
-Simulation results are provided in Section V.
-Finally, we conclude the paper in Section VI.
+In Section V, we study the MLE model in the joint device activity and data detection context.
+Simulation results are provided in Section VI.
+Finally, we conclude the paper in Section VII.
 In this paper, lower-case, boldface lower-case, and boldface upper-case letters denote scalars, vectors, and matrices, respectively.
@@
 We use to denote the binary variable with for active and for inactive devices.
-We consider the case where the devices are in the near-field region, i.e., the distance between the devices and the BS is less than the Rayleigh distance, which is given by where and denote the antenna aperture and the wavelength, respectively [14].
-We assume that both the deterministic LoS channel and statistical multipath channels exist, and the channel between device and the BS can be modeled as the following Rician fading model [14, Eq.
-(133)]:
-where is the LoS channel between device and the BS, and
+We consider the case where some devices are located in the near-field region of the BS, and others in the far-field region, as shown in Fig.
+1.
+The radius of the near-field region is indicated by the Rayleigh distance, which is given by where and denote the antenna aperture and the wavelength, respectively [15].
+Notice when the carrier frequency is set as GHz, and the number of antennas is set as and the corresponding Rayleigh distances are approximately m, m, and m, respectively.
+Consequently, the number of near-field devices can be large due to the substantial Rayleigh distance.
+We assume that both the deterministic LoS channel and statistical multipath channels exist, and the channel between device and the BS can be modeled as the following Rician fading model:
+where is the LoS channel between device and the BS. For a near-field device the mean vector can be expressed as [15, Eq. (133)]:
 where denotes the channel gain between device and the BS, is the distance between the device and the -th antenna, and is the distance between the device and the midpoint of the antenna array.
-The correlation matrix is given by [14, Eq.
+Furthermore, the correlation matrix is given by [15, Eq.
 (134)]:
 where is the number of scatterers between device and the BS, is the intensity attenuation caused by the -th scatterer, denotes the channel gain between the -th scatterer and the device and is the channel vector between the -th scatterer and the BS with a similar structure to (2), i.e., where and denote the corresponding channel gain and distances.
-Notice that and thus is a low-rank matrix when the number of scatterers
+Notice that and thus is a low-rank matrix when the number of scatterers For a far-field device the vector can be represented by the conventional plane wave model, and the correlation matrix can be modeled as where is the large-scale fading coefficient [44].
 Each device is preassigned a unique signature sequence with length for activity detection.
 During the uplink pilot stage, all active devices synchronously transmit their signature sequences to the BS as random access requests.
-Following this, the received signal at the BS can be expressed as
+Following this, the received signal at the BS can be expressed as [6, 7, 8]
 where is the additive i.i.d. white Gaussian noise with variance Denote as the activity indicator vector.
@@
 To make the problem more computationally tractable, we relax the constraint to as in [7, 8, 9], and obtain the following relaxed problem:
+When solving the MLE problem (10) for device activity detection, we only need to know the mean and the correlation matrix of each device’s channel, without requiring the knowledge of the channel vector i.e., the channel state.
+Notice that appears in (4) because of the system model [6, 7, 8], and appears in (10) due to the probability distribution function of the complex Gaussian distribution [57].
 In this paper, we focus on the algorithm design and performance analysis for the MLE problem (10).
-Notice that the dimension of the covariance matrix in problem (10) is which is significantly different from the case of uncorrelated channels.
-In the latter case, the covariance matrix involved in the MLE formulation is [5, 6, 7].
-This difference makes problem (10) much more challenging to solve.
-III Proposed CD Algorithms for Device Activity Detection
+Notice that the dimension of the covariance matrix in problem (10) is which is much larger than that in the case of uncorrelated channels [6, 7, 8], i.e., Therefore, the complexity of updating and storing the matrix in the CD algorithm is much higher than that in the case of uncorrelated channels.
+In addition, in the case of uncorrelated channels, updating a single coordinate of corresponds to a rank- update of the covariance matrix, which enables a closed-form solution for each coordinate update [6, 7, 8].
+However, in the correlated case, due to (II-B), the rank- update property does not hold, making it impossible to obtain a closed-form solution for coordinate updates.
+These differences make problem (10) much more challenging to solve.
+III Efficient CD Algorithms for Device Activity Detection
 In this section, we focus on designing efficient algorithms for solving the MLE problem (10).
-Notice that problem (10) is nonconvex due to the concavity of We propose two algorithms for solving the MLE problem: an exact CD and an inexact CD.
-In the exact CD algorithm, we solve the one-dimensional optimization subproblem exactly, while in the inexact CD algorithm, we solve it inexactly in order to reduce the computational cost.
+Notice that problem (10) is nonconvex due to the concavity of We first present the exact CD algorithm for solving the MLE problem, which solves the one-dimensional optimization subproblem exactly.
+In contrast, the proposed inexact CD algorithm solves this subproblem approximately, thereby achieving improved numerical stability and reduced computational cost.
 Furthermore, we study a special case that allows us to further reduce the complexity of both the exact CD and inexact CD algorithms by utilizing the low-rank property of the near-field channel correlation matrix.
@@
 where and are the derivatives of and respectively.
-Notice that the left-hand side in (24) is a polynomial of degree so problem (24) can be solved using a polynomial root-finding algorithm [56, Chap.
+Notice that the left-hand side in (24) is a polynomial of degree so problem (24) can be solved using a polynomial root-finding algorithm [58, Chap.
 6].
@@
 The procedure of solving problem (19) exactly is outlined in lines 6 and 7 of Algorithm 1.
-The above proposed method is suitable to solve problem (19) for small due to the following reasons.
-When is large, significant numerical errors arise in two steps: (i) the eigenvalue decomposition of i.e., the computation of and satisfying (20), and (ii) the root-finding of the polynomial of degree in (24).
-The error in step (i) makes the coefficients of and inaccurate, while the error in step (ii) makes it impossible to obtain a solution to problem (23).
+The above method is only suitable to solve problem (19) for small As observed in Section VI-A, the exact CD algorithm may diverge when is large.
+The observed divergence can be attributed to two main factors.
+First, for large the structures of and may cause their values to become very close to zero during certain coordinate updates.
+As a result, even small numerical errors in computing and can lead to significant inaccuracies in the objective function value of problem (23).
+Second, the root-finding process in problem (24) can produce numerically inaccurate solutions when the leading coefficient of the polynomial becomes excessively large, a scenario that becomes increasingly likely as increases.
 To address these issues, we propose a method for solving problem (19) inexactly, as detailed below.
@@
 Theorem 1 demonstrates that, at each coordinate update, although the inexact CD algorithm does not solve problem (19) exactly, it finds a good approximate solution that sufficiently decreases the objective function, especially when is large.
-This guarantees the convergence of the inexact CD algorithm (to a stationary point).
+Furthermore, building upon the techniques introduced in [59], this theorem also establishes the convergence of the inexact CD algorithm to a stationary point.
 III-A3 Complexity
@@
 In massive MIMO systems, updating and storing dominate the complexity of both the exact and inexact CD algorithms.
-As follows, we consider a special case where the complexity of updating and storing can be significantly reduced.
+In the next subsection, we consider a special case where the complexity of updating and storing can be significantly reduced.
+III-A4 Differences between Presented Algorithms and Those in [9]
+First, we emphasize that the optimization problem (10) is significantly different from problem (7) in [9].
+This difference is primarily due to two key factors: (i) the dimension of the covariance matrix is significantly larger than that in [9]; and (ii) the rank-1 update property, which holds for the covariance matrices in [9], does not apply to
+Next, we compare the exact CD algorithm with the CD algorithm in [9].
+Both of them transform the optimization problem into a polynomial root-finding problem, but unlike [9] where coefficients are computed analytically, this paper proposes applying eigenvalue decomposition to determine the polynomial coefficients.
+Finally, we compare the inexact CD algorithm with the inexact CD algorithm in [9].
+The main difference between the two algorithms is in their handling of the parameter In [9], the parameter is dynamically adjusted at each coordinate update.
+However, in this paper, the parameter is fixed to a sufficiently large positive number to ensure convergence.
+This difference arises because in [9], the computational complexity of the objective function is implying that adjusting does not incur any additional computational cost.
+In contrast, in this paper, adjusting would result in a computational complexity for the objective function of at least Consequently, we fix to reduce the complexity of coordinate updates.
 III-B A Special Case Study
-In this subsection, we consider a special case in which the proposed exact CD and inexact CD can be accelerated.
+In this subsection, we consider a special case in which the presented exact CD and inexact CD can be accelerated.
 The motivation arises from the fact that within a cell, there may be only a few devices in the near-field region of the BS, whereas the majority of devices are located in the far-field region.
@@
 We further assume that the summation of the correlation matrices of the first devices is rank-deficient, i.e.,
-Notice that Then, a sufficient condition for Assumption 1 to hold is that there are few near-field devices (i.e., is small) and few scatterers in the near-field channels (i.e., is small).
+Notice that Then, the limitation to Assumption 1 is twofold: first, the number of devices whose channels are correlated needs to be sufficiently smaller than the number of antennas; second, their channel correlation matrices must be low-rank.
+For example, consider a scenario where the number of antennas at the BS is There are potential devices in the cell, but only devices are in the near-field region, and the channels of the other devices are uncorrelated.
+Assuming the number of scatterers in the near-field channel is (i.e., ), we then have
 The following proposition demonstrates that under Assumption 1, () can be transformed into a matrix where only the upper-left submatrix is nonzero.
@@
 Case II:
-For each device , the channel follows where is the large-scale fading coefficient.
+For each device the channel follows where is the large-scale fading coefficient.
 We assume that the set of signature sequences is the same in both Case I and Case II, with each sequence being randomly generated from a specific distribution, such as the uniform distribution over the discrete set as in [9].
 IV-A Signal Statistical Dimension
-First we define the signal statistical dimensions and in Case I and Case II respectively:
+The main idea of this subsection is to define and compare signal statistical dimensions in Case I and Case II.
+We infer that a larger signal statistical dimension corresponds to better detection performance because the received signal can be transformed by a unitary transformation.
+In the transformed signal vector, only the covariance matrix of a subvector provides information about The length of this subvector is the signal statistical dimension.
+We then compare the upper bounds of the signal statistical dimensions in Case I and Case II, which indicates the relationship between the system parameters and the signal statistical dimensions.
+We now define the signal statistical dimensions and in Case I and Case II respectively:
 The statistical dimensions and are closely related to the detection performance in Case I and Case II due to the following reason.
@@
 IV-B Identifiability of True Activity Vector
-The phase transition analysis in [7] demonstrates that, a crucial condition for the MLE model (10) to perform effectively is that the true activity indicator vector needs to be uniquely identifiable.
-This implies that there should be no other vector such that We now investigate the identifiability of in both Case I and Case II.
+The main idea of this subsection arises from the phase transition analysis in [8], which demonstrates that a crucial condition for the MLE model (10) to perform effectively is that the true activity indicator vector needs to be uniquely identifiable.
+This implies that there should be no other vector such that We demonstrate that the unique identifiability of is more likely to be achieved in Case I than in Case II.
+Furthermore, inspired by the unique identifiability of we infer that the mutual interference between different devices is weaker in Case I compared to Case II.
+We now investigate the identifiability of in both Case I and Case II.
 Since both and are multivariate Gaussian distributions, their equivalence implies that their corresponding means and covariances must be identical.
@@
 where follows from the Cauchy-Schwartz inequality and equality holds if and only if there exists a constant such that We observe from (54) that and are closer to orthogonal than and implying that the mutual interference between devices and in Case I is lower than that in Case II.
-The above analysis demonstrates that the channel correlation among different antennas can reduce the correlation between devices and thereby improve the detection performance.
+The above analysis demonstrates that the channel correlation among different antennas can reduce the mutual interference between devices and thereby improve the detection performance.
 Therefore, when the signal statistical dimensions in Case I and Case II are equal (i.e., is large), the MLE model (10) exhibits superior detection performance in Case I (than in Case II).
-V Simulation Results
+V Joint Device Activity and Data Detection
+This section aims to show that the MLE model (10) can be applied to the joint device activity and data detection problem, in which each device is associated with multiple signature sequences and can embed a few information bits [10].
+Suppose that each active device intends to send bits of information to the BS, where is a small integer.
+To perform joint activity and data detection, we assign each device a unique signature sequence set of size :
+where When device is active and needs to send bits of data, it selects a single sequence from its assigned set and sends this sequence to the BS.
+The BS then detects which sequences have been transmitted to perform the device activity detection and data decoding.
+The process of performing joint activity and data detection using the covariance-based approach is as follows.
+Let indicate whether sequence of device (i.e., ) is transmitted.
+Since at most one sequence is selected by each device, satisfies where indicates that device is active, and indicates that device is inactive.
+Then, the received signal at the BS, as well as its vectorization, can be expressed as follows:
+Therefore, we have where the mean is
+and the covariance matrix is given by
+Using the same techniques presented in Section II-B, the MLE problem can be expressed as
+To make the problem more computationally tractable, we relax the binary constraint to as in [7, 8, 9], and we relax the combination constraint as in [10]. Then we obtain the following relaxed problem:
+Therefore, by solving problem (60) using the exact CD and inexact CD algorithms presented in this paper, we can effectively perform joint activity and data detection.
+VI Simulation Results
 We consider an mMTC system with a single cell of radius m, where all potential devices are uniformly distributed.
@@
 The signature sequences are generated randomly and uniformly from the discrete set as in [9].
-V-A Computational Efficiency of Proposed Algorithms
-In this subsection, we showcase the efficiency of the proposed exact and inexact CD algorithms for solving the device activity detection problem (10).
+Notice that there are enough devices in the near-field region because, as mentioned in Section II-A, the Rayleigh distance can be sufficiently large, especially when the number of antennas is large.
+All experiments are implemented in Matlab 2021b on a computer equipped with an Intel Core i7-11700 CPU and 32 GB of memory.
+This is for a fair comparison in computational complexity with existing algorithms.
+In practical implementations, more efficient programming languages such as C/C++ would be adopted, and the code would also be optimized for even more efficient execution.
+VI-A Computational Performance of Presented Algorithms
+In this subsection, we demonstrate the computational performance of the exact and inexact CD algorithms for solving the device activity detection problem (10).
 In the inexact CD algorithm, we set the parameter to
+Table I shows the convergence behavior of the exact CD algorithm for different ’s.
+For each we record the proportion of experiments in which the exact CD algorithm converges out of experiments.
+We observe that for the system parameter settings listed in Table I, the exact CD algorithm always converges when (corresponding to ).
+Conversely, the algorithm always diverges when (for most cases ).
+Comparing Setting 1 with Setting 2, we observe that as the number of antennas decreases from to the convergence threshold for increases from to Similarly, comparing Setting 1 and Setting 3, we find that reducing the signature sequence length from to also increases the convergence threshold for from to Furthermore, by comparing Setting 1 with Settings 4 and 5, we observe that parameters and do not influence the convergence threshold of
 In Fig.
-1, we show the decrease of the objective function values of the proposed exact and inexact CD algorithms with respect to the running time with and The outputs of the exact and inexact CD algorithms are denoted as and respectively.
+2, we show the decrease of the objective function values of the exact and inexact CD algorithms with respect to the running time with and The outputs of the exact and inexact CD algorithms are denoted as and respectively.
 We then define as the minimum of and We observe from Fig.
-1(a) that when the objective function values of exact and inexact CD can gradually decrease and converge to the same value.
+2 that when and the objective function values of exact and inexact CD can gradually decrease and converge to the same value.
 Moreover, exact CD is significantly more efficient than inexact CD.
@@
 In Fig.
-1(b), we observe that when the objective function value of the inexact CD algorithm still converges, but that of the exact CD algorithm cannot decrease and diverges.
-This occurs because when is large (corresponding to a large ), the exact CD algorithm fails to solve the one-dimensional subproblem (19) due to the numerical errors associated with the high-dimensional eigenvalue decomposition and polynomial root-finding.
+2(d), we observe that when the objective function value of the inexact CD algorithm still converges, but that of the exact CD algorithm cannot decrease and diverges.
 This observation shows that exact CD is more efficient when is small, while inexact CD is more robust when is large.
 In Fig.
-2, we assume that Assumption 1 holds true and plot the decrease of the objective function values of the proposed exact and inexact CD algorithms as well as their accelerated versions by Algorithm 2 with respect to the running time with and In Fig.
-2(a), we set in Assumption 1, and for (implying that ).
+3, we assume that Assumption 1 holds true and plot the decrease of the objective function values of the exact and inexact CD algorithms as well as their accelerated versions by Algorithm 2 with respect to the running time with and In Fig.
+3(a), we set in Assumption 1, and for (implying that ).
 Consequently, we have in (30).
 Notice that in this case, there are devices with channel correlation matrix Fig.
-2(a) shows that, in this scenario, the objective function value of the accelerated inexact CD converges to and it is more computationally efficient than the inexact CD algorithm.
+3(a) shows that, in this scenario, the objective function value of the accelerated inexact CD converges to and it is more computationally efficient than the inexact CD algorithm.
 However, neither exact CD nor its accelerated version converges.
 This is because neither exact CD nor its accelerated version can solve the one-dimensional optimization subproblem (19) when the channel correlation matrix of the device To demonstrate that the exact CD algorithm can also be accelerated when it converges, we remove devices with channel correlation matrix Therefore, in Fig.
-2(b), we set in Assumption 1, and for all We observe from Fig.
-2(b) that the accelerated version of exact CD converges, and it is more efficient than exact CD.
+3(b), we set in Assumption 1, and for all We observe from Fig.
+3(b) that the accelerated version of exact CD converges, and it is more efficient than exact CD.
 The same observation applies to the inexact CD algorithm as well.
-A summary of the proposed CD algorithms is as follows.
+A summary of the presented CD algorithms is as follows.
 First, the exact CD algorithm is more computationally efficient than the inexact CD algorithm when the rank of the channel correlation matrix is small.
@@
 Due to its robustness in solving the subproblem, we adopt the inexact CD algorithm to solve problem (10) in different scenarios in the following two subsections.
-V-B Numerical Validation of Detection Performance Analysis
+VI-B Numerical Validation of Detection Performance Analysis
 In this subsection, we verify the detection performance analysis in Section IV.
 We consider the device activity detection problem under Case I and Case II and perform MLE model (10) in both cases.
-In order to demonstrate the impact of channel correlation matrices on the detection performance as compared to the traditional uncorrelated Rayleigh fading channels, we set for all for a fair comparison.
+Notice that the key difference between Case I and Case II lies in the channel correlation matrices.
+In order to highlight the impact of the channel correlation matrices on detection performance while eliminating the effect of channel means, we set the channel means to zero in this subsection.
+The termination condition of the inexact CD algorithm is set as
 In Fig.
-3, we show the detection performance for different values of with and The “probability of error” in the figure is defined as the point where the probability of missed detection (PM) and probability of false alarm (PF) are equal.
+4, we show the detection performance for different values of with and The “probability of error” in the figure is defined as the point where the probability of missed detection (PM) and probability of false alarm (PF) are equal.
 We observe that increasing can improve the detection performance in both cases.
@@
 This also verifies when is larger, the rank of is larger, and thus the signal statistical dimension is larger, and the detection performance is better.
+The crossover for and is because when is small, the signal statistical dimension is the same for both cases, as it cannot exceed When is large, the larger the is, the larger the rank of will be, and consequently, the larger the signal statistical dimension.
+Therefore, the detection performance for is better than that for
 Fig.
@@
 This is due to when is large, the signal statistical dimensions are equal in both cases; however, in Case I, the true activity vector is more easily identifiable compared to Case II, as discussed in Section IV-B.
-Additionally, we also observe that in Case I, a smaller tends to result in better detection performance for larger This can be explained by the fact that when is smaller, the cosine similarity in (54) is likely to be lower, indicating a weaker correlation between different devices, which in turn leads to improved detection performance.
+Additionally, we also observe that in Case I, a smaller tends to result in better detection performance for larger This can be explained by the fact that when is smaller, the cosine similarity in (54) is likely to be lower, indicating a weaker interference between different devices, which in turn leads to improved detection performance.
+Fig.
+6 plots the detection performance for various values of with parameters and Notice that the sparsity level of active devices ranges from to We observe that the detection performance in both cases deteriorates as increases, exhibiting a similar trend to that as in Fig.
+5 when increases.
+When is large, the detection performance in Case I is superior to that in Case II.
+Additionally, within Case I, a smaller value corresponds to better detection performance for large This is because, as discussed in Section IV-B, the interference between devices in Case I is less than that in Case II.
+Moreover, in Case I, a smaller value corresponds to a smaller cosine similarity in (54), resulting in less interference between devices.
 Fig.
 7 shows the detection performance for various values of with and We observe that as increases, the detection performance in Case I with values of and significantly outperforms that of Case I with and Case II.
-This suggests that when is large and the rank of is large in Case I, the detection performance is superior compared to the case when the rank of is low in Case I, and it is also superior to Case II.
-V-C Detection Performance Comparison with Benchmarks
+The detection performance of Case I with is worse than that of Case I with and when is large, because the signal statistical dimension in Case I with is lower than that in Case I with and The detection performance of Case II is worse than that of Case I with and because, in Case I, the identifiability of the true activity vector is easier to hold, and there is less mutual interference between different devices.
+VI-C Detection Performance Comparison with Benchmarks
 In this subsection, we compare the detection performance of the proposed MLE-based approach with the following benchmarks:
@@
 OAMP is an improved version of AMP, which is applicable to more types of signature sequences.
-In the scenario of Rician fading channels, i.e., we ignore the LoS channel and assume the prior distribution to be in both methods.
+Notice that while AMP and OAMP can be applied under correlated Rayleigh fading channels, they cannot be directly applied to correlated Rician fading.
+The reason is that the denoiser and state evolution in these algorithms assume zero channel means, which cannot be directly extended to scenarios with nonzero channel means.
+In the scenario of Rician fading channels, i.e., we ignore the mean and assume the prior distribution to be in both methods.
 Notice that both methods require the knowledge of the probability that a device is active, i.e., but the MLE does not require this prior information.
 Mismatched MLE: We apply the vanilla MLE-based approach in [44]. Specifically, we use to replace the channel correlation matrix in the MLE formulation.
+The termination condition of the inexact CD algorithm is set as and we allow other benchmarks to run sufficiently long until they converge.
+Notice that AMP [32] and OAMP [35] are CS-based methods, while the MLE proposed in this paper and mismatched MLE are covariance-based methods.
+The key difference between the CS-based method and the covariance-based method is that the former jointly estimates the channels and device activities, whereas the latter estimates only the device activities.
+Other covariance-based methods include SPICE [30], which can be applied to the Rayleigh fading scenario but encounters technical difficulties in the Rician fading scenario.
 In Fig.
-6, we show the detection performance under correlated Rician/Rayleigh fading channels for various values of with and It can be observed from Fig.
+8, we show the detection performance under correlated Rician/Rayleigh fading channels for various values of with and It can be observed from Fig.
 8 that increasing can improve the detection performance of all compared methods.
 In Fig.
-6(a), we observe that the detection performance of MLE is significantly better than that of all the other benchmarks.
+8(a), we observe that the detection performance of MLE is significantly better than that of all the other benchmarks.
 The main reason for this is that these benchmarks use inaccurate channel distributions: both AMP and OAMP do not utilize the LoS channel, and the mismatched MLE roughly approximates In Fig.
-6(b), we observe that the detection performance of MLE is slightly better than that of AMP and OAMP when is large and significantly better than that of mismatched MLE.
+8(b), we observe that the detection performance of MLE is slightly better than that of AMP and OAMP when is large and significantly better than that of mismatched MLE.
 In this case, all methods except mismatched MLE use accurate channel distributions.
@@
 In Fig.
-7(a), due to the use of inaccurate channel distributions, the detection performance of AMP, OAMP, and mismatched MLE is significantly worse than that of MLE.
+9(a), due to the use of inaccurate channel distributions, the detection performance of AMP, OAMP, and mismatched MLE is significantly worse than that of MLE.
 Moreover, when is large, the detection performance of mismatched MLE becomes better than that of AMP and OAMP.
 In Fig.
-7(b), we observe that the detection performances of MLE, AMP, and OAMP are very close, and the detection performance of mismatched MLE is significantly worse than theirs.
+9(b), we observe that the detection performances of MLE, AMP, and OAMP are very close, and the detection performance of mismatched MLE is significantly worse than theirs.
 Fig.
@@
 Fig.
-8(a) illustrates that in the Rician fading scenario, the detection performance of MLE is significantly better than that of the other methods.
+10(a) illustrates that in the Rician fading scenario, the detection performance of MLE is significantly better than that of the other methods.
 Fig.
-8(b) shows that in the Rayleigh fading scenario, the detection performance of MLE, AMP, and OAMP is very close, with the mismatched MLE performing worse than these methods.
+10(b) shows that in the Rayleigh fading scenario, the detection performance of MLE, AMP, and OAMP is very close, with the mismatched MLE performing worse than these methods.
 Moreover, MLE outperforms AMP and OAMP when is large.
-VI Conclusion
+VI-D Joint Device Activity and Data Detection
+In this subsection, we study the joint device activity and data detection problem and demonstrate the detection performance of the MLE model (60).
+Fig.
+11 plots the detection performance for various values of with and We observe from Fig.
+11 that when is small, the smaller the number of information bits the better the detection performance.
+In addition, as increases, for bit and the detection performance of the MLE model improves and gradually becomes equal.
+Note that when there is no data transmission, and only activity detection is performed.
+This indicates that in the joint device activity and data detection problem, transmitting bits of information (when is small) does not significantly deteriorate the detection performance of the MLE model, especially when is large.
+VII Conclusion
 This paper investigates the device activity detection problem for NFC, where the near-field channel can be modeled as correlated Rician fading.
 The corresponding device activity detection problem is formulated as an MLE problem.
-Two efficient algorithms for solving the MLE problem, i.e., exact and inexact CD, are proposed.
+We present the exact CD algorithm and propose the inexact CD algorithm with improved numerical stability for solving the MLE problem.
 In addition, an analysis of the detection performance of MLE is provided.
-Simulation results demonstrate the efficiency of the proposed algorithms and verify the correctness of the detection performance analysis.
+Simulation results demonstrate the efficiency of the presented algorithms and verify the correctness of the detection performance analysis.
 Appendix A Proof of Theorem 1
@@
 Since is continuous and is in the bounded region is Lipschitz continuous. That is, there exists a constant such that for any the following inequality holds:
-Furthermore, we have a quadratic upper bound on [57, Proposition A.24]:
+Furthermore, we have a quadratic upper bound on [60, Proposition A.24]:
 Setting and in (64), we obtain
@@
 where is derived from
-and the properties of the projection operator Similar techniques are used in the proof of [8, Proposition 4].
+and the properties of the projection operator Similar techniques are used in the proof of [9, Proposition 4].
 Substituting into (69) and combining it with (A), we obtain (29).